OpenStudy (anonymous):

sin^2x-cos^2x=0 solve for the equation for the interval [0, 2pi)

OpenStudy (anonymous):

Therefore: \[\sin^2(x) = \cos^2(x)\]

OpenStudy (anonymous):

\(\large sin^2x-cos^2x=0 \) \(\large (sinx+cosx)(sinx-cosx)=0 \) can u do it from here?

OpenStudy (anonymous):

my choices all have pi/4 in it not sure how to get there please help

OpenStudy (anonymous):

just set each factor equal to zero and solve.... for example: sinx + cosx = 0 can you solve this?

OpenStudy (anonymous):

if we solve dpalnc's equations, we have: \[\sin(x) = \cos(x)\] or: \[\sin(x) = -\cos(x)\]

OpenStudy (anonymous):

for sinx = cosx, what angle give the same sine and cosine??? hint: there are two of 'em.... in [0, 2pi)

OpenStudy (anonymous):

\[\frac{ \pi }{ 4 }, \frac{ 3\pi }{4 }, \frac{ 5\pi }{ 4 },\frac{ 7\pi }{ 4 }\]so the answer would be

OpenStudy (anonymous):

@dpaInc though this does not relate to the question, how did you make your avatar?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):